OpenStudy (anonymous):

plz prove the answer: 1.d/dx(logb u)=1/u(logb exponent e) du/dx 2.d/dx(ln u)=1/u du/dx i need the solution plz...

zepdrix (zepdrix):

\[\Large \frac{d}{dx}\log_bu=\frac{1}{u}\left(\log_b ?\right)\frac{du}{dx}\] What's going on on the right side in that log? I don't quite understand.